OpenStudy (anonymous):

Is this statement true or false? log2 45 – log2 3 = log2 5 + log2 3

OpenStudy (anonymous):

it's true :) because 45/3=5*3

OpenStudy (anonymous):

\[ \log_2(45)-\log_23=\\ \log_2(3^2\times 5)-\log_23= \\2\log_23+\log_25-\log_23= \\\log_23+\log_25 \]